A 800 kg machine is mounted on four identical springs of total spring constant k = 1.5 x 106 N/m and negligible damping. The machine is subjected to a harmonic external force of amplitude Fo = 490 N and frequency of 180 rpm. The amplitude of the motion of the machine
1
Expert's answer
2020-10-02T14:51:22-0400

"A_0=\frac{F_0}{m\sqrt{(\omega_0^2-\omega^2)^2+4\beta^2\omega^2}}"


"180(rpm)=3(c^{-1})"


"\omega=2\pi\nu=2\cdot3.14\cdot3=18.84(rad/s)"


"\omega_0^2=k/m=1.5\cdot10^6/800=1875(rad^2/s^2)"


For negligible damping "A_0\approx\frac{F_0}{m\sqrt{(\omega_0^2-\omega^2)^2}}=\frac{F_0}{m(\omega_0^2-\omega^2)}\to"


"A_0=\frac{F_0}{m(\omega_0^2-\omega^2)}=\frac{490}{800\cdot(1875-18.84^2)}=4\cdot10^{-4}(m)"